# Statistical modeling to adjust for time trends in adaptive platform trials utilizing non-concurrent controls

### v1

@@ 4.2 Results
Setting 2A. In Figures 7 and 8 the type I error rate and power of the model with the calendar time adjustment is compared to the period adjustment and the separate analysis under varying unit size and $\lambda$, respectively. Corresponding results for other experimental arms are presented in Figures S6 and S7 in the supplementary material. For Figure 7, unit sizes in the range from 25 to 750 (see Table 1) were considered. The type I error rate control for different time unit lengths is dependent on the time trend pattern. In case of linear and inverted-U trend, the type I error rate is maintained for moderately sized calendar time units, and slightly inflated for units larger than 600 patients. Under the seasonal time trend, we observe deflation of the type I error for unit sizes $\geq 200$. In case of the stepwise trend, the type I error rate control is only given for very small unit sizes ($<50$ patients). The loss of type I error rate control for larger calendar unit sizes is due to the fact that larger unit sizes lead to greater pooling of the concurrent and non-concurrent data. In trials with sudden changes in the time trend (such as the considered stepwise trend pattern), the type I error inflation for naive pooling is more pronounced, therefore the control is only given for smaller sizes. The type I error rate would be controlled in settings with stepwise trend if treatment arms were only allowed to enter the trial at the beginning of a calendar time unit. This, however, was not the case in our simulations. Depending on the unit size, the model with calendar time adjustment can lead to power improvements as compared to the period adjustment, especially in the setting with inverted-U or linear time trend. Nevertheless, as the type I error rate is not maintained in all cases, the choice of the interval length and the resulting trade-off between type I and type II errors needs to be carefully assessed.
@@
Figure 8: Setting 2A: Type I error rate and power of the regression model with calendar time adjustment compared to the regression model with period adjustment and separate analysis with respect to the strength of the time trend $\lambda$ under different time trend patterns. Here, $n=250$ in each experimental arm, $N_{p}=750$ in case of inverted-U trend (corresponding approximately to the middle of the trial), and calendar time unit size of 100 are used. Treatment arm 3 is being evaluated.
@@
Figure 9: Setting 2B: Type I error rate and power of the mixed model with period and calendar time adjustments as uncorrelated and autocorrelated random effects, compared to the fixed effect regression model with period adjustment with respect to the pattern and strength of the time trend. Sample sizes of $n=250$ in each experimental arm and $N_{p}=750$ in case of inverted-U trend (corresponding approximately to the middle of the trial) are assumed. In case of calendar time adjustment, unit size of 100 patients is considered. Treatment arm 3 is being evaluated.
@@
Figure 10: Setting 3: Type I error rate and power of the mixed model with interaction and period or calendar time adjustments, compared to the fixed effect regression model with period adjustment with respect to the time trend case and its strength. Sample sizes of $n=250$ in each experimental arm are assumed. In case of calendar time adjustment, unit size of 100 patients is considered. Treatment arm 3 is being evaluated.
4.3 Implementation
The simulation study was performed using the NCC R package [17], which contains functions implementing all considered analysis approaches, as well as functions for data simulation and performing simulation studies [18].
@@ 5 Case Study
We exemplify the use of the proposed methods on data from the ABBV-8E12 trial [19] for Progressive Supranuclear Palsy (PSP), a rare neurodegenerative disorder affecting balance, body movements, vision, and speech, eventually resulting in death. The ABBV-8E12 study evaluated the safety and efficacy of different doses (2000mg - here denoted as arm 1; and 4000mg - denoted as arm 2) of the experimental treatment Tilavonemab in a randomized parallel group trial, using placebo as a common control group. The trial recruited patients over a period of approximately 2 years, from January 2017 to March 2019. When revisiting this data, we consider the Progressive Supranuclear Palsy Rating Scale (PSPRS) as an endpoint, a sum score summing up the values for 28 different items, which measure the progression of the disease across different domains (lower values of this score indicate better outcomes) [20]. To demonstrate the potential impact of time trends under the null hypothesis, for illustrative purposes we compare the baseline measurements of the sum score between treatment and control groups (instead of the change from baseline as in the original trial). For the baseline measurements, we know that due to randomization there is no treatment effect compared to placebo for both arms, such that all observed treatment effects are either due to bias or sampling error. We furthermore demonstrate the use of the proposed models on this data.
To emulate a two-period platform trial, we modified the original multi-arm trial data and assumed that the second experimental arm (4000mg of Tilavonemab) enters the ongoing trial at a later time point. Therefore, we set the cut-off for the periods (i.e., the time point where treatment arm 2 enters the platform) to 1.6.2018 and exclude observations from patients allocated to arm 2 prior to this cut-off date from the analysis. The resulting dataset consists of 315 patients, with sample sizes of 123, 126, and 66 for the placebo arm, and treatment arms 1 and 2, respectively. For approaches that adjust for calendar time, we use a unit length of 90 days, which divides the trial into 9 calendar time units, each of them corresponding to approximately 3 months. Note that as the patient recruitment was not uniform, the numbers of patients in each unit are different, ranging from 4 in the first unit to 96 in the 8th. Figure 11 visualizes the considered data, showing the division into periods (red dashed line) and calendar time units (grey dashed lines).
@@
Subsequently, we analyze the dataset using the proposed modeling approaches, as well as the separate and pooled analyses, and compare the estimated treatment effects, their standard errors, and p-values. In the case of the mixed models, we only consider the calendar time adjustment, as the analyzed trial only consists of two periods, which would lead to a very unstable estimation of the random effect in case of the period adjustment. We focus on evaluating the efficacy of the 2nd experimental arm against control, since this arm has non-concurrent control data available. Table 2 summarizes the obtained results. Since the considered responses were measured at the baseline date in a randomized trial, we can assume that the null hypothesis holds, thus the underlying treatment effect is 0 for both arms. We evaluate this null hypothesis ($H_{0,2}:\theta_{2}=0$) against the two-sided alternative ($H_{1,2}:\theta_{2}\neq 0$). None of the analysis approaches reject the null hypothesis. The most accurate estimate of the treatment effect (-0.0123) is provided by the fixed effect model adjusting for periods, followed by the fixed effect model with calendar time adjustment (-0.0682). These models also result in the highest p-values against the null hypothesis (p=0.995 and p=0.971, respectively). The second closest estimate is provided by the spline regression, resulting in estimated effects of 0.1795 and 0.6239 for the period and calendar time adjustment, respectively. In both model classes (fixed effect models and spline regressions), the period adjustment leads to a better effect estimate than the calendar time adjustment. The standard errors of the estimates are comparable for all considered models, ranging from 1.8223 (mixed model with autocorrelated random effects and calendar time adjustment) to 1.9002 (fixed effect model with calendar time adjustment). We can also observe that all modeling approaches yield estimates that are closer to the true effect than the pooled analysis, which introduces a considerable bias to the effect estimation (-1.7030) and also provides the lowest p-value (p=0.336).
In the ABBV-8E12 trial, the primary endpoint was the change from baseline in the PSPRS score after 52 weeks. We also investigated potential time trends for this endpoint, but observed no substantial drift in this outcome during the trial (data not shown). Hence, this data suggests that time trends would not pose a problem in this particular trial, and the fact that patients with a milder stage of PSP were recruited later in the trial is accounted for by adjusting for the baseline response. For an assessment of the efficiency of other choices for the primary endpoint see Yousefi et al. [21].
@@
Fixed effect model Periods -0.0123 1.8961 0.995
Calendar time units -0.0682 1.9002 0.971
Mixed model Calendar time units -0.9028 1.8326 0.623
Mixed model (AR1) Calendar time units -0.9730 1.8223 0.594
Spline regression Periods 0.1795 1.8851 0.924
Calendar time units 0.6239 1.8794 0.740
Pooled analysis - -1.7030 1.7660 0.336
Separate analysis - -0.4104 2.0323 0.840
Table 2: Treatment effect estimates, standard errors and two-sided p-values obtained when comparing the second treatment arm against the control using the proposed approaches.
@@ 6 Discussion
We evaluated the performance of the proposed methods in terms of the type I error rate and statistical power in a simulation study, considering a wide range of scenarios. We specified under which conditions the fixed effect model adjusting for periods leads to power gains as compared to the separate analysis. Moreover, we showed that the calendar time adjustment in fixed effect models can, in some scenarios, lead to power gains as compared to the period adjustment while still controlling the type I error rate. In the considered scenarios, the mixed models without interaction were not suitable for evaluating the efficacy of late-entering treatment arms in the presence of time trends, as they do not preserve the type I error rate. Lee and Wason [8] and Bofill Roig et al. [9] investigated a linear regression model that includes time as a continuous, rather than categorical variable. It was shown that this model only controls the type I error rate if the underlying time trend has a linear pattern. The spline regression can be viewed as a generalization of this model, fitting one polynomial function for each period or calendar time interval, rather than one linear function for the whole trial, to adjust for time. We demonstrated via simulations that modeling time via a spline function controls the type I error rate for all considered scenarios with continuous time trends. Furthermore, it was shown that the models with fixed factor period that also include a term for interaction between treatment and time as a random effect, can improve the robustness and decrease the inflation of the type I error rate compared to the model without this interaction, if the time trends are not equal across arms. The proposed analysis approaches focus only on trials with continuous endpoints. The extension of the presented methods to other types of endpoints, such as binary or time-to-event endpoints, remains open for future research.
@@
Non-concurrent controls share some characteristics with historical controls, as they both refer to data collected prior to the data on the treatment under study, and thus both might introduce bias in the estimates and inflate the type I error rate if included in the analysis [22, 23]. However, compared to historical controls, non-concurrent controls comprise patients who have been part of the same trial framework as the patients allocated to the investigated treatment. Therefore, they typically share the inclusion and exclusion criteria and endpoint assessment procedures [24]. Methods for incorporating historical controls in randomized clinical trials have been widely discussed in the last years, and could be analogously used in the context of platform trials to incorporate NCC data [25]. Frequentist methods include for instance the “test-then-pool” approach, where either separate or pooled analysis is performed based on a frequentist test assessing the equality of the CC and historical data [15]; dynamic pooling, which assigns a weight parameter to the historical data to control their proportion in the analysis [26]; or propensity score methods that can be used to adjust for imbalances in baseline covariates between historical and concurrent controls [27]. Bayesian approaches encompass the power prior method, which down-weights the historical information by introducing a power parameter to their likelihood function [28]; commensurate power prior, an extension to the power prior approach that directly parameterizes the commensurability of the historical and CC data [29]; and meta-analytic-predictive (MAP) prior approaches, which uses the historical information to derive a MAP prior distribution for the CC data [30, 31]. Other Bayesian methods, such as dynamic borrowing, were also proposed for incorporating external data into the trial analysis in order to obtain more information for the decision-making [32, 33]. However, in contrast to the modeling approaches to incorporate NCC data, these methods do not directly adjust for the time trend, but rather down-weight the historical information in order to ameliorate the bias due to heterogeneity in the data. Hence, they still introduce some bias if time trends are present. On the contrary, the model-based approaches for incorporating NCC lead to unbiased estimates provided that the model assumptions are met. A detailed overview of methods discussed in the context of historical, non-concurrent, and/or external controls is provided in the scoping review by Bofill Roig et al. [25].
@@
While analysis using concurrent controls only is still considered the preferred approach, under certain circumstances, such as trials for rare diseases, the use of non-concurrent controls can be valuable [11, 25]. In such cases, methods for addressing the potential bias due to time trends should be employed [11]. The proposed models in this work are in line with these suggestions, increasing the precision of the treatment effect estimators, while mitigating the biases that could arise due to temporal changes.
@@
Supplementary Material Formal definition of the time variables, further specifications of the models, and additional results from the simulation study can be found in the pdf file with supplementary material. The GitHub repository (https://github.com/pavlakrotka/NCC_FreqModels) contains the R code to reproduce the results presented in this paper. Authors’ Contributions M.B.R., M.P. and P.K. conceived this research. G.H. and M.G. established data access and harmonized data across datasets. P.K. prepared the initial draft of the text and performed the simulations and statistical analyses. All authors discussed the results, provided comments and reviewed the manuscript.

### v2

@@ 4.2 Results
Setting 2A. In Figures 8 and 9 the type I error rate and power of the model with the calendar time adjustment is compared to the period adjustment and the separate analysis under varying unit size and $\lambda$, respectively. Corresponding results for other experimental arms are presented in Figures S8 and S9 in the supplementary material. For Figure 8, unit sizes in the range from 25 to 750 (see Table 1) were considered. The length for which the type I error rate is still controlled depends on the assumed time trend pattern and its strength. In case of moderately strong linear and inverted-U trends, the type I error rate is maintained for calendar time units of approximately 1/4 of the length of the considered trial, and slightly inflated for units larger than 400 patients. Under the seasonal time trend, we observe deflation of the type I error for unit sizes larger than 200. In case of the stepwise trend, the type I error rate control is only given for very small unit sizes (<50<50 patients).
In the case considered here, with calendar time units chosen such that treatment arms can enter or leave within the interval, adjusting for these intervals leads to pooling of trial data with different randomization scheme, since the number of active arms changes. The loss of type I error rate control, which can be observed especially for larger calendar unit sizes, is then due to the greater effect of data pooling. In trials with sudden changes in the time trend (such as the considered stepwise trend pattern), the type I error inflation for naive pooling is even more pronounced.
Depending on the unit size, the model with calendar time adjustment can lead to power improvements as compared to the period adjustment. In particular, in trials where the patient response does not change substantially over time (i.e. trials with linear or inverted-U time trend pattern and small $\lambda$ relative to the treatment effect), power gains can be achieved by considering calendar time units larger than periods, while still maintaining the type I error rate control (see Figure 8).
As the choice of the length of the calendar time interval determines the resulting trade-off between type I and type II error rates and needs to be carefully assessed in simulations when planning the platform trial. If the experimental arms would be allowed to enter and leave the trial only at the beginning of a calendar time unit, the type I error rate control would be ensured regardless of the time trend pattern and strength, as every change in the randomization design would be adjusted for. However, under this setting, the model with calendar time adjustment does not offer any gain in power as compared to the period adjustment. We did not consider this design in our simulations, as we aimed to evaluate the general case with arbitrary unit sizes.
@@
Figure 9: Setting 2A: Type I error rate and power of the regression model with calendar time adjustment compared to the regression model with period adjustment and separate analysis with respect to the strength of the time trend $\lambda$ under different time trend patterns. Here, $n=250$ in each experimental arm, $N_{p}=750$ in case of inverted-U trend (corresponding approximately to the middle of the trial), and calendar time unit size of 100 are used. Treatment arm 3 is being evaluated. In the power plots for settings with inverted-U, linear, and seasonal time trends, the lines for the regression models with period (red) and calendar time (blue) adjustment are overlapping.
@@
Figure 10: Setting 2B: Type I error rate and power of the mixed model with period and calendar time adjustments as uncorrelated and autocorrelated random effects, compared to the fixed effect regression model with period adjustment with respect to the pattern and strength of the time trend. Sample sizes of $n=250$ in each experimental arm and $N_{p}=750$ in case of inverted-U trend (corresponding approximately to the middle of the trial) are assumed. In case of calendar time adjustment, unit size of 100 patients is considered. Treatment arm 3 is being evaluated. In settings with inverted-U, linear, and seasonal time trends, the lines for the mixed effects models with autocorrelated random effects and period (brown) and calendar time (magenta) adjustment overlap. In the setting with stepwise time trend, the lines for the mixed effect models with period adjustment and uncorrelated (blue) and autocorrelated (brown) random effects overlap.
@@
As shown in [9], the type I error rate is controlled for different time trends across arms when an interaction term is included as a fixed effect. However, such a model offers no power gains compared to the separate analysis.
Figure 11: Setting 3: Type I error rate and power of the mixed model with interaction and period or calendar time adjustments, compared to the fixed effect regression model with period adjustment with respect to the time trend case and its strength. Sample sizes of $n=250$ in each experimental arm are assumed. In case of calendar time adjustment, unit size of 100 patients is considered. Treatment arm 3 is being evaluated. In the power plot under equal time trends, the lines for all model-based approaches (red, blue, and yellow) overlap.
4.3 Summary of main results
Below we summarize the advantages and limitations of the considered methods and the main results from the simulation study and describe the influence of certain design parameters.
Fixed effect regression with period adjustment: Fixed effect models with period adjustment maintain the type I error rate and lead to a power increase compared to the separate analysis in case of equal time trends between arms that are additive on the model scale, for arbitrary time trend patterns. However, a certain overlap in time between the experimental treatment arms is required in order to achieve a power gain. The amount of the power increase depends on the size of the NCC data and the amount of overlap.
Spline regression with period adjustment: Spline regression maintains the type I error rate in scenarios where the time trend pattern is sufficiently smooth and equal across all arms. Moreover, it can lead to a power increase as compared to the fixed effect model with period adjustment, especially in cases with little to no overlap between the experimental treatment arms.
Mixed effect regression with period adjustment: Mixed models where the period is modeled as random effect do not control the type I error rate in the presence of time trends. The inflation is more pronounced in cases with moderate time trends compared to cases with strong time trends. Mixed models with a fixed effect period and a random effect for the interaction between treatment and period control the type I error rate in trials with equal time trends, and reduce the inflation in settings with unequal time trends, but do not eliminate it completely. In cases with unequal time trends, the type I error rate is controlled by a regression model that includes an interaction between treatment and time as a fixed effect [9]. This method, however, offers no power gains compared to the separate analysis.
Period vs. calendar time adjustment: Additionally to the period adjustment investigated in previous works, we considered to adjust for calendar time units that are, unlike the periods, independent of the trial progress. Using this adjustment, the number of factor levels in the resulting regression model, or number of inner knots for the spline function, can be controlled by the additional design parameter $c_{length}$ (calendar time unit length). Depending on the pattern and strength of the time trend, adjusting for larger units than those given by the periods can leads to power gains, while still controlling the type I error rate. Type I error rate control would be ensured in cases where new experimental arms only enter and leave the trial at the beginning of a calendar time unit, but this setting would not lead to substantial power gains as compared to the period adjustment. Also note that the last calendar time interval is cut at the time point where the evaluated arm leaves the platform to avoid including future data to the analysis, which might not be available by the time of assessing the given arm.
Overlap between experimental arms: We showed that certain overlap between the experimental arms is necessary for the fixed effect regression model with period adjustment to achieve power gains as compared to the separate analysis. This is, however, not required by the spline regression, which uses continuous, rather than categorical time adjustment and leads to power increase even in cases with no overlap, i.e. if only one experimental arm is active at any given time.
@@ 5 Case Study
We exemplify the use of the proposed methods in a case study based on a real-world clinical trial. It revisits real data from a trial for progressive supranuclear palsy [19] and analyzes the baseline measurements, hence illustrates the use of the modeling approaches in a setting where the null hypothesis is assumed to hold.
Additionally, in the supplementary material (Sect. D), we present a second case study that mimics the design of the FLAIR platform trial for chronic lymphocytic leukemia [20, 21, 22] using simulated data. It incorporates certain design elements from the FLAIR trial and demonstrates the application of methods in a scenario involving two efficacious treatment options.
We reanalyze data from the ABBV-8E12 trial [19] for progressive supranuclear palsy (PSP), a rare neurodegenerative disorder affecting balance, body movements, vision, and speech, eventually resulting in death. The ABBV-8E12 study evaluated the safety and efficacy of different doses (2000mg - here denoted as arm 1; and 4000mg - denoted as arm 2) of the experimental treatment Tilavonemab in a randomized parallel group trial, using placebo as a common control group. The trial recruited patients over a period of approximately 2 years, from January 2017 to March 2019. When revisiting this data, we consider the Progressive Supranuclear Palsy Rating Scale (PSPRS) as an endpoint, a sum score summing up the values for 28 different items, which measure the progression of the disease across different domains (lower values of this score indicate better outcomes) [23]. To demonstrate the potential impact of time trends under the null hypothesis, for illustrative purposes we compare the baseline measurements of the sum score between treatment and control groups (instead of the change from baseline as in the original trial). For the baseline measurements, we know that due to randomization there is no treatment effect compared to placebo for both arms, such that all observed treatment effects are either due to bias or sampling error. We furthermore demonstrate the use of the proposed models on this data.
To emulate a two-period platform trial, we modified the original multi-arm trial data and assumed that the second experimental arm (4000mg of Tilavonemab) enters the ongoing trial at a later time point. Therefore, we set the cut-off for the periods (i.e., the time point where treatment arm 2 enters the platform) to 1.6.2018 and exclude observations from patients allocated to arm 2 prior to this cut-off date from the analysis. The resulting dataset consists of 315 patients, with sample sizes of 123, 126, and 66 for the placebo arm, and treatment arms 1 and 2, respectively. The sample sizes per arm and period are presented in Figure 12. For approaches that adjust for calendar time, we use a unit length of 90 days, which divides the trial into 9 calendar time units, each of them corresponding to approximately 3 months. Note that as the patient recruitment was not uniform, the numbers of patients in each unit are different, ranging from 4 in the first unit to 96 in the 8th. Figure 13 visualizes the considered data, showing the division into periods (red dashed line) and calendar time units (grey dashed lines).
@@
Subsequently, we analyze the dataset using the proposed modeling approaches, as well as the separate and pooled analyses, and compare the estimated treatment effects, their standard errors, and p-values. In the case of the mixed models, we only consider the calendar time adjustment, as the analyzed trial only consists of two periods, which would lead to a very unstable estimation of the random effect in case of the period adjustment. We focus on evaluating the efficacy of the 2nd experimental arm against control, since this arm has non-concurrent control data available. Table 2 summarizes the obtained results. Since the considered responses were measured at the baseline date in a randomized trial, we can assume that the null hypothesis holds, thus the underlying treatment effect is 0 for both arms. We evaluate this null hypothesis ($H_{0,2}:\theta_{2}=0$) against the two-sided alternative ($H_{1,2}:\theta_{2}\neq 0$). None of the analysis approaches reject the null hypothesis. The most accurate estimate of the treatment effect (-0.012) is provided by the fixed effect model adjusting for periods, followed by the fixed effect model with calendar time adjustment (-0.068). These models also result in the highest p-values against the null hypothesis (p=0.995 and p=0.971, respectively). The second closest estimate is provided by the spline regression, resulting in estimated effects of 0.178 and 0.624 for the period and calendar time adjustment, respectively. In both model classes (fixed effect models and spline regressions), the period adjustment leads to a better effect estimate than the calendar time adjustment. The standard errors of the estimates are comparable for all considered models, ranging from 1.822 (mixed model with autocorrelated random effects and calendar time adjustment) to 1.900 (fixed effect model with calendar time adjustment). We can also observe that all modeling approaches yield estimates that are closer to the true effect than the pooled analysis, which introduces a considerable bias to the effect estimation (-1.7030) and also provides the lowest p-value (p=0.336).
In the ABBV-8E12 trial, the primary endpoint was the change from baseline in the PSPRS score after 52 weeks. We also investigated potential time trends for this endpoint, but observed no substantial drift in this outcome during the trial (data not shown). Hence, this data suggests that time trends would not pose a problem in this particular trial, and the fact that patients with a milder stage of PSP were recruited later in the trial is accounted for by adjusting for the baseline response. For an assessment of the efficiency of other choices for the primary endpoint, see [24].
Figure 12: Scheme of the platform trial considered in the case study for progressive supranuclear palsy (ABBV-8E12 trial). Sample sizes per arm and period are shown.
@@
Fixed effect model Periods -0.012 1.896 0.995
Calendar time units -0.068 1.900 0.971
Mixed model Calendar time units -0.903 1.833 0.623
Mixed model (AR1) Calendar time units -0.973 1.822 0.594
Spline regression Periods 0.178 1.885 0.924
Calendar time units 0.624 1.879 0.740
Pooled analysis - -1.703 1.766 0.336
Separate analysis - -0.410 2.032 0.840
Table 2: ABBV-8E12 trial: Treatment effect estimates, standard errors and two-sided p-values obtained when comparing the second treatment arm against the control using the proposed approaches.
@@ 6 Discussion
We evaluated the performance of the proposed methods in terms of the type I error rate and statistical power in a simulation study, considering a wide range of scenarios. We specified under which conditions the fixed effect model adjusting for periods leads to power gains as compared to the separate analysis. Moreover, we showed that the calendar time adjustment in fixed effect models can, in some scenarios, lead to power gains as compared to the period adjustment while still controlling the type I error rate. In the considered scenarios, the mixed models without interaction were not suitable for evaluating the efficacy of late-entering treatment arms in the presence of time trends, as they do not preserve the type I error rate. [8] and [9] investigated a linear regression model that includes time as a continuous, rather than categorical variable. It was shown that this model only controls the type I error rate if the underlying time trend has a linear pattern. The spline regression can be viewed as a generalization of this model, fitting one polynomial function for each period or calendar time interval, rather than one linear function for the whole trial, to adjust for time. We demonstrated via simulations that modeling time via a spline function controls the type I error rate for all considered scenarios with sufficiently smooth time trends. Another method that uses smoothing across time to estimate the treatment effect is the Bayesian Time Machine [10]. However, the performance of the Time Machine strongly depends on the choice of prior distributions. Moreover, it requires the specification of the calendar time intervals. The spline regression with period adjustment, where the placement of the knots is given by the periods, on the other hand, only requires the choice of the degree of the polynomial function, with cubic splines being a suitable default option, as they ensure type I error rate control even in cases with complex time trend patterns. Furthermore, it was shown that the models with fixed factor period that also include a term for the interaction between treatment and time as a random effect, can improve the robustness and decrease the inflation of the type I error rate compared to the model without this interaction, if the time trends are not equal across arms. The proposed analysis approaches focus only on trials with continuous endpoints. The extension of the presented methods to other types of endpoints, such as binary or time-to-event endpoints, remains open for future research.
@@
Non-concurrent controls share some characteristics with historical controls, as they both refer to data collected prior to the data on the treatment under study, and thus both might introduce bias in the estimates and inflate the type I error rate if included in the analysis [25, 26]. However, compared to historical controls, non-concurrent controls comprise patients who have been part of the same trial framework as the patients allocated to the investigated treatment. Therefore, they typically share the inclusion and exclusion criteria and endpoint assessment procedures [27]. Methods for incorporating historical controls in randomized clinical trials have been widely discussed in the last years, and could be analogously used in the context of platform trials to incorporate NCC data [28]. Frequentist methods include for instance the “test-then-pool” approach, where either separate or pooled analysis is performed based on a frequentist test assessing the equality of the CC and historical data [15]; dynamic pooling, which assigns a weight parameter to the historical data to control their proportion in the analysis [29]; or propensity score methods that can be used to adjust for imbalances in baseline covariates between historical and concurrent controls [30]. Bayesian approaches encompass the power prior method, which down-weights the historical information by introducing a power parameter to their likelihood function [31]; commensurate power prior, an extension to the power prior approach that directly parameterizes the commensurability of the historical and CC data [32]; and meta-analytic-predictive (MAP) prior approaches, which uses the historical information to derive a MAP prior distribution for the CC data [33, 34]. Other Bayesian methods, such as dynamic borrowing, were also proposed for incorporating external data into the trial analysis in order to obtain more information for the decision-making [35, 36]. However, in contrast to the modeling approaches to incorporate NCC data, these methods do not directly adjust for the time trend, but rather down-weight the historical information in order to ameliorate the bias due to heterogeneity in the data. Hence, they still introduce some bias if time trends are present. On the contrary, the model-based approaches for incorporating NCC lead to unbiased estimates provided that the model assumptions are met. A detailed overview of methods discussed in the context of historical, non-concurrent, and/or external controls is provided in the scoping review by [28].
Recently, [37] considered the use of NCCs in trials with varying eligibility criteria across arms. In this case, also the use of concurrent control subjects that did not meet the eligibility criteria of the corresponding treatment arm may introduce bias in treatment control comparisons. They refer to non-concurrent controls for a given experimental arm as patients who do not belong to the same concurrently randomized cohort. The concurrently randomized cohort comprises only patients that were eligible for the same set of treatments and were randomized using the same allocation method [37]. In this paper we assumed that for all arms the same eligibility criteria apply. If this is not the case, one can fit the proposed models in the subgroup of patients that satisfy the eligibility criteria of the tested treatment arm.
In addition to the risk of bias due to time trends, there might be other sources of bias when utilizing non-concurrent controls. As platform trials are designed to run over multiple years, results from completed arms could be published before the whole trial ends. This may influence the specification of the analysis methods and thus the trial integrity for arms that are still active or are being considered to enter the platform [1]. For instance, it may affect the willingness of the intervention owners to include new treatment arms in the platform, making platform trials with a control with a random low in the outcome more attractive to join. Conversely, a platform trial with a random high can be a deterrent for the intervention owners.
@@
While analysis using concurrent controls only is still considered the preferred approach, under certain circumstances, such as trials for rare diseases, the use of non-concurrent controls can be valuable [11, 28]. In such cases, methods for addressing the potential bias due to time trends should be employed [11]. The proposed models in this work aim to address these time trends and can improve the precision of the treatment effect estimators, while mitigating the biases that could arise due to temporal changes. However, given that the performance of the analysis methods depends on many factors, the implications of including non-concurrent controls and the behavior of the considered methods must be investigated on a case-by-case basis when planning a specific platform trial.
Supplementary Material Formal definition of the time variables, further specifications of the models, and additional results from the simulation study can be found in the pdf file with supplementary material.
The GitHub repository (https://github.com/pavlakrotka/NCC_FreqModels) contains the R code to reproduce the results presented in this paper.
@@
Authors’ Contributions M.B.R., M.P. and P.K. conceived this research. G.H. and M.G. established data access and harmonized data across datasets. P.K. prepared the initial draft of the text and performed the simulations and statistical analyses. All authors discussed the results, provided comments and reviewed the manuscript.